A bicycle and rider have a combined mass of 114.0 kg. How many calories of heat are generated in the brakes when the bicycle comes to a stop from a speed of 32.2 km/h? kcal
Answer #1

Here,

initial speed , u = 32.2 km/hr

u = 32.2 * 1000/3600 m/s = 8.94 m/s

mass = 114 Kg

heat generated = loss in kinetic energy

heat generated = 0.50 * 114 * 8.94^2

heat generated = 4560 J

heat generated = 4560/4186 Kcal

heat generated = 1.09 Kcal

the heat generated is 1.09 Kcal

> Where did the 4186 come from

james white Tue, Apr 12, 2022 9:38 AM